# Towards Trustworthy Reranking: A Simple yet Effective Abstention Mechanism

### v4

@@ 4.1 Models and Datasets
An extensive benchmark is built to evaluate the abstention mechanisms presented in Section 3. We collect six open-source reranking datasets Lhoest et al. (2021) in three different languages, English, French and Chinese: stackoverflowdupquestions-reranking Zhang et al. (2015), denoted StackOverflow in the experiments, askubuntudupquestions-reranking Lei et al. (2015), denoted AskUbuntu, scidocs-reranking Cohan et al. (2020), denoted SciDocs, mteb-fr-reranking-alloprof-s2p Lefebvre-Brossard et al. (2023), denoted Alloprof, CMedQAv1-reranking Zhang et al. (2017), denoted CMedQAv1, and Mmarco-reranking Bonifacio et al. (2021), denoted Mmarco.
We also collect 22 open-source models for evaluation on each of the six datasets. These include models of various sizes, bi-encoders and cross-encoders, monolingual and multilingual: ember-v1 (paper to be released soon), llm-embedder Zhang et al. (2023), bge-base-en-v1.5, bge-reranker-base/large Xiao et al. (2023), multilingual-e5-small/large, e5-small-v2/large Wang et al. (2022), msmarco-MiniLM-L6-cos-v5, msmarco-distilbert-dot-v5, ms-marco-TinyBERT-L-2-v2, ms-marco-MiniLM-L-6-v2 Reimers and Gurevych (2021), stsb-TinyBERT-L-4, stsb-distilroberta-base, multi-qa-distilbert-cos-v1, multi-qa-MiniLM-L6-cos-v1, all-MiniLM-L6-v2, all-distilroberta-v1, all-mpnet-base-v2, quora-distilroberta-base and qnli-distilroberta-base Reimers and Gurevych (2019) [^7] We take the liberty of evaluating models on languages in which they have not been trained, as we posit it is possible to detect relevant patterns in the relevance scores in any case. We discuss this assertion in Section 5.2..
@@ 4.2 Instance-Wise Metrics
In this work, we rely on three of the metrics most commonly used in the IR setting (e.g., in the MTEB leaderboard Muennighoff et al. (2022)): Average Precision (AP), Normalized Discounted Cumulative Gain (NDCG) and Reciprocal Rank (RR): AP Zhu (2004) computes a proxy of the area under the precision-recall curve; NDCG Järvelin and Kekäläinen (2002) measures the relevance of the top-ranked items by discounting those further down the list using a logarithmic factor; and RR Voorhees et al. (1999) computes the inverse rank of the first relevant element in the predicted ranking. For all three metrics, higher values indicate better performance. Note that, when relevant, we rely on their averaged versions across multiple instances: mAP, mNDCG and mRR.
@@ 4.3 Assessing Abstention Performance
The compromise between abstention rate and performance in predictive modeling represents a delicate balance that requires careful consideration El-Yaniv and Wiener (2010). First, we want to make sure that an increasing abstention rate implies increasing performance, otherwise the mechanism employed is ineffective or even deleterious. Secondly, we aim to abstain in the best possible way for every abstention rate, i.e., we want to maximize the growth of the performance-abstention curve.
Inspired by the risk-coverage curve El-Yaniv and Wiener (2010); Geifman and El-Yaniv (2017), we evaluate abstention strategies by reporting the normalized Area Under the performance-abstention Curve, where the x-axis corresponds to the abstention rate of the strategy and the y-axis to the achieved average performance regarding metric function $m$. Our evaluation protocol can be divided into four steps, assuming access to a test set, $\mathcal{S}^{\prime}$.
1. Multi-thresholding. First of all, we evaluate abstention mechanism $\rho_{\bm{\theta}}$ (Equation 5) on test set $\mathcal{S}^{\prime}$ for a list of abstention thresholds $\tau_{1}<...<\tau_{n}\in\mathbb{R}$. For a given $\tau\in\mathbb{R}$, the performance of mechanism $\rho_{\bm{\theta}}$ on $\mathcal{S}^{\prime}$ and its corresponding abstention rate are respectively denoted by
@@
$\mathcal{S}_{\tau}^{\prime}$ being the set of predicted instances, i.e., $\mathcal{S}_{\tau}^{\prime}=\left\{\left(x,\mathbf{y}\right)\in\mathcal{S}^{\prime}\ |\ \rho_{\bm{\theta}}\left(x,\tau\right)\neq\perp\right\}$. We collect $n$ abstention rates, $\left(R_{\tau_{i}}\left(\rho_{\bm{\theta}}\right)\right)_{i=1}^{n}$, and associated performances, $\left(P_{\tau_{i},m}\left(\rho_{\bm{\theta}}\right)\right)_{i=1}^{n}$, and compute the area under the curve $AUC_{m}\left(\rho_{\bm{\theta}}\right)$ for metric function $m$.
2. Random mechanism evaluation. Second, we evaluate $\tilde{\rho}$, the mechanism that performs random (i.e., ineffective) abstention. Intuitively, random abstention shows flat performance as the abstention rate increases, equal to $P_{-\infty,m}\left(\rho_{\bm{\theta}}\right)$ (Equation 7).
3. Oracle evaluation. Then, we evaluate the oracle $\rho^{*}$ on the same task. Intuitively, the oracle has access to the test labels and can therefore select the best instances for a given abstention rate. $AUC_{m}\left(\rho^{*}\right)$ thus upper-bounds $AUC_{m}\left(\rho_{\bm{\theta}}\right)$.
4. Normalization. We finally compute normalized AUC. Formally,
@@ 5.1 Abstention Performance
Figure 2: Performance-abstention curves for the ember-v1 model on the StackOverflow dataset. All methods show increasing curves, indicating effective abstention. The reference-based method, $u_{\text{lin}}$, stands out above the others, demonstrating superior abstention performance. Note that $u^{*}$ and $\tilde{u}$ respectively denote the oracle and the random abstention confidence functions.
Table 1: Abstention performance assessment. nAUCs (in %) averaged model-wise, for each method, dataset and metric.
Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$
Dataset Metric
SciDocs mAP 49.2 58.7 4.7 65.4
mNDCG 54.9 62.9 10.2 66.6
mRR 80.2 80.3 41.3 80.5
AskUbuntu mAP 24.1 16.2 7.3 22.0
mNDCG 28.1 15.7 8.6 24.5
mRR 32.7 16.6 14.0 26.2
StackOverflow mAP 20.7 24.6 35.7 42.6
mNDCG 21.0 25.0 35.6 42.6
mRR 21.4 24.8 36.1 42.8
Alloprof mAP 16.6 17.0 19.3 29.7
mNDCG 16.8 16.9 18.8 29.3
mRR 16.9 16.7 19.1 28.9
CMedQAv1 mAP 24.4 23.3 22.3 30.6
mNDCG 25.0 23.5 22.3 30.7
mRR 26.5 24.2 23.4 31.5
Mmarco mAP 30.1 31.0 39.4 34.0
mNDCG 30.5 30.9 38.8 34.1
mRR 30.1 31.0 39.6 34.3
Average mAP 27.5 28.5 21.5 37.4
mNDCG 29.4 29.1 22.4 38.0
mRR 34.6 32.3 28.9 40.7
We report abstention nAUC (Equation 8) for all methods described in Section 3, averaged model-wise (Table 1), and illustrate the dynamics of the performance-abstention curves for the ember-v1 model on the StackOverflow dataset in Figure 2.
@@
Reference-based abstention works better. The value of reference-based abstention approaches is demonstrated in Table 1, in which we notice that the linear-regression-based method $u_{\text{lin}}$ performs on average largely better than all reference-free baselines, edging out the best baseline strategy ($u_{\text{std}}$) by almost 10 points on the mAP metric.
Having shown the superiority of reference-based abstention methods across the entire benchmark, we conduct (unless otherwise specified) the following experiments on the ember-v1 model using the mAP as a metric function, and compare $u_{\text{lin}}$ to $u_{\text{std}}$, the best reference-free method mAP-wise.
@@
Figure 3: Abstention performance (nAUC in %) vs. no-abstention performance (mAP), for all models and datasets, using $u_{\text{lin}}$ as a confidence function. Each data point represents a model-dataset pair.
Table 2: Pearson’s correlations between no-abstention mAP and abstention nAUC, for each dataset.
Dataset Correlation
SciDocs 0.85
AskUbuntu 0.38
StackOverflow 0.90
Alloprof 0.56
CMedQAv1 0.94
Mmarco 0.63
General 0.73
We investigate whether a correlation exists between abstention effectiveness and model raw performance (i.e., without abstention) on the reranking task. For each model-dataset pair, we compute both no-abstention mAP and abstention nAUC, using $u_{\text{lin}}$ as a confidence function (Figure 3, Table 2).
A better ranker implies better abstention. Figure 3 and Table 2 showcase a clear positive correlation between model raw performance and abstention effectiveness, albeit with disparities from one dataset to another. Logically, better reranking methods exhibit more calibrated score distributions, enabling better abstention performance. This finding reinforces the interest of implementing abstention mechanisms in high-performance retrieval systems.
@@ 5.3 Threshold Calibration
Table 3: Threshold calibration analysis. MAEs between target and achieved abstention rates and performance levels (in %) (model: ember-v1, dataset: StackOverflow, metric: mAP).
Target Abstention 10% 50% 90%
Rate $u_{\text{std}}$ 1.10 1.80 1.08
$u_{\text{lin}}$ 1.07 1.79 1.09
mAP $u_{\text{std}}$ 1.24 1.52 3.09
$u_{\text{lin}}$ 1.22 1.33 1.51
In practical industrial settings, a major challenge with abstention methods lies in choosing the right threshold to ensure a desired abstention rate or performance level on new data. To evaluate threshold calibration quality, we sequentially choose three target abstention rates (10%, 50%, and 90%) and rely on the reference set [^8] In this setting, access to a reference set is essential in order to determine which threshold value corresponds to a given abstention rate or level of performance. to infer the corresponding threshold and level of performance (here, the mAP). Then, the Mean Absolute Error (MAE) between target and achieved values on test instances is reported (Table 3), averaging across 1000 random seeds on the StackOverflow dataset [^9] For example, if the mAP corresponding to a target abstention rate of 50% in the reference set is 0.6 and the values achieved on test set on 2 different seeds are 0.5 and 0.7, the MAE will be equal to $1/2\times(|0.6-0.5|+|0.6-0.7|)=0.1$..
$u_{\text{lin}}$ is better calibrated than $u_{\text{std}}$. The lower part of Table 3 shows that the MAE between the target and achieved mAP increases with the abstention rate, which is logical because a higher abstention rate reduces the number of evaluated instances and therefore increases volatility. However, we observe that the MAE increases much less significantly for the $u_{\text{lin}}$ method than for the $u_{\text{std}}$ method, with the latter having a MAE twice as high at a target abstention rate of 90%. This provides strong evidence that $u_{\text{lin}}$ is more reliable at high abstention rates.
@@ 5.4 Domain Adaptation Study
Table 4: Domain adaptation study. nAUCs (in %) averaged test-set-wise (model: ember-v1, metric: mAP).
Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$
Reference set
Alloprof 30.1 32.7 17.1 31.4
AskUbuntu 28.5 32.1 19.8 34.2
CMedQAv1 27.9 31.1 17.2 34.2
Mmarco 27.5 32.2 18.9 34.9
SciDocs 23.0 23.5 22.3 11.6
StackOverflow 29.3 30.7 11.9 20.6
Average 27.7 30.4 17.9 27.8
In a practical setting, it is rare to have a reference set that perfectly matches the distribution of samples seen at test time. To measure the robustness of our method to strong data drifts, we fit them on a given dataset and evaluate on all others.
Reference-based approaches are globally not robust to domain changes. From Table 4, we see on average that the reference-based method underperforms reference-free baselines when fitted on the "wrong" reference set. However, certain reference datasets seem to enable methods to generalize better overall, even out of distribution, which is mainly an artifact of the number of positive documents per instance (further insights in Appendix C).
@@ 5.5 Reference Set Minimal Size
Figure 4: Reference set size analysis. nAUC vs. reference set size for $u_{\text{lin}}$ and $u_{\text{std}}$ on all datasets (model: ember-v1, metric: mAP).
Table 5: Reference set size analysis. Minimum reference set size required for $u_{\text{lin}}$ to outperform $u_{\text{std}}$ (model: ember-v1, metric: mAP).
Dataset Break-Even
SciDocs 12
AskUbuntu 9
StackOverflow 19
Alloprof 58
CMedQAv1 50
Mmarco 80
Average 38
@@
Small reference sets suffice. From Figure 4 and Table 5, we note that $u_{\text{lin}}$ requires only a fairly small amount of reference data (less than 40 samples on average) to outperform $u_{\text{std}}$, which demonstrates the effectiveness of using a reference set even when data is scarce. This demonstrates the applicability of our method to practical industrial contexts, the labeling of around 50 samples appearing perfectly tractable in such settings.

### v5

@@ 4.1 Models and Datasets
An extensive benchmark is built to evaluate the abstention mechanisms presented in Section 3. We collect six open-source reranking datasets (Lhoest et al., 2021) in three different languages, English, French and Chinese: stackoverflowdupquestions-reranking (Zhang et al., 2015), denoted StackOverflow in the experiments, askubuntudupquestions-reranking (Lei et al., 2015), denoted AskUbuntu, scidocs-reranking (Cohan et al., 2020), denoted SciDocs, mteb-fr-reranking-alloprof-s2p (Lefebvre-Brossard et al., 2023), denoted Alloprof, CMedQAv1-reranking (Zhang et al., 2017), denoted CMedQAv1, and Mmarco-reranking (Bonifacio et al., 2021), denoted Mmarco.
We also collect 22 open-source models for evaluation on each of the six datasets. These include models of various sizes, bi-encoders and cross-encoders, monolingual and multilingual: ember-v1 (paper to be released soon), llm-embedder (Zhang et al., 2023), bge-base-en-v1.5, bge-reranker-base/large (Xiao et al., 2023), multilingual-e5-small/large, e5-small-v2/large-v2 (Wang et al., 2022), msmarco-MiniLM-L6-cos-v5, msmarco-distilbert-dot-v5, ms-marco-TinyBERT-L-2-v2, ms-marco-MiniLM-L-6-v2 (Reimers & Gurevych, 2021), stsb-TinyBERT-L-4, stsb-distilroberta-base, multi-qa-distilbert-cos-v1, multi-qa-MiniLM-L6-cos-v1, all-MiniLM-L6-v2, all-distilroberta-v1, all-mpnet-base-v2, quora-distilroberta-base and qnli-distilroberta-base (Reimers & Gurevych, 2019). [^7] We take the liberty of evaluating models on languages in which they have not been trained, as we posit it is possible to detect relevant patterns in the relevance scores in any case. We discuss this assertion in Section 5.2.
@@ 4.2 Instance-Wise Metrics
In this work, we rely on three of the metrics most commonly used in the IR setting (e.g., MTEB leaderboard (Muennighoff et al., 2022)): Average Precision (AP), Normalized Discounted Cumulative Gain (NDCG) and Reciprocal Rank (RR): AP (Zhu, 2004) computes a proxy of the area under the precision-recall curve; NDCG (Järvelin & Kekäläinen, 2002) measures the relevance of the top-ranked items by discounting those further down the list using a logarithmic factor; and RR (Voorhees et al., 1999) computes the inverse rank of the first relevant element in the predicted ranking. For all three metrics, higher values indicate better performance. When relevant, we rely on their averaged versions across multiple instances: mAP, mNDCG and mRR.
@@ 4.3 Assessing Abstention Performance
The compromise between abstention rate and performance in predictive modeling represents a delicate balance that requires careful consideration (El-Yaniv & Wiener, 2010). First, we want to make sure that an increasing abstention rate implies increasing performance, otherwise the mechanism employed is ineffective or even deleterious. Secondly, we aim to abstain in the best possible way for every abstention rate, i.e., we want to maximize the growth of the performance-abstention curve.
Inspired by the risk-coverage curve (El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017), we evaluate abstention strategies by reporting the normalized Area Under the performance-abstention Curve, where the x-axis corresponds to the abstention rate of the strategy and the y-axis to the achieved average performance regarding metric function $m$. Our evaluation protocol is as follows, assuming access to a test set, $\mathcal{S}^{\prime}$:
  1. 1.
    Multi-thresholding. First of all, we evaluate abstention mechanism $\rho_{\bm{\theta}}$ (Equation 5) on test set $\mathcal{S}^{\prime}$ for a list of abstention thresholds $\tau_{1}<...<\tau_{n}\in\mathbb{R}$. For a given $\tau\in\mathbb{R}$, the performance of mechanism $\rho_{\bm{\theta}}$ on $\mathcal{S}^{\prime}$ and its corresponding abstention rate are respectively denoted by
@@
    $\mathcal{S}_{\tau}^{\prime}$ being the set of predicted instances, i.e., $\mathcal{S}_{\tau}^{\prime}=\left\{\left(x,\mathbf{y}\right)\in\mathcal{S}^{\prime}\ |\ \rho_{\bm{\theta}}\left(x,\tau\right)\neq\perp\right\}$. We collect $n$ abstention rates, $\left(R_{\tau_{i}}\left(\rho_{\bm{\theta}}\right)\right)_{i=1}^{n}$, and associated performances, $\left(P_{\tau_{i},m}\left(\rho_{\bm{\theta}}\right)\right)_{i=1}^{n}$, and then compute the area under the performance-abstention curve $AUC_{m}\left(\rho_{\bm{\theta}}\right)$ for metric function $m$.
  2. 2.
    Random baseline. Second, we evaluate $\tilde{\rho}$, the mechanism that performs random (i.e., ineffective) abstention. Intuitively, random abstention shows flat performance as the abstention rate increases, equal to $P_{-\infty,m}\left(\rho_{\bm{\theta}}\right)$ (Equation 7). [^8] A random (ineffective) abstention method cannot differentiate between good and bad instances. Thus, in expectation, random abstention on the test dataset results in a constant performance whatever the abstention rate.
  3. 3.
    Oracle evaluation. Then, we evaluate the oracle $\rho^{*}$ on the same task. Intuitively, the oracle has access to the test labels and can therefore select the best instances for a given abstention rate. $AUC_{m}\left(\rho^{*}\right)$ thus upper-bounds $AUC_{m}\left(\rho_{\bm{\theta}}\right)$.
  4. 4.
    Normalization. We finally compute normalized AUC. Formally,
@@ 5.1 Abstention Performance
[…]
tableAbstention performance. nAUCs in % averaged model-wise, for each method, dataset and metric. Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ Dataset Metric SciDocs mAP 49.2 58.7 4.7 65.4 mNDCG 54.9 62.9 10.2 66.6 mRR 80.2 80.3 41.3 80.5 AskUbuntu mAP 24.1 16.2 7.3 22.0 mNDCG 28.1 15.7 8.6 24.5 mRR 32.7 16.6 14.0 26.2 StackOverflow mAP 20.7 24.6 35.7 42.6 mNDCG 21.0 25.0 35.6 42.6 mRR 21.4 24.8 36.1 42.8 Alloprof mAP 16.6 17.0 19.3 29.7 mNDCG 16.8 16.9 18.8 29.3 mRR 16.9 16.7 19.1 28.9 CMedQAv1 mAP 24.4 23.3 22.3 30.6 mNDCG 25.0 23.5 22.3 30.7 mRR 26.5 24.2 23.4 31.5 Mmarco mAP 30.1 31.0 39.4 34.0 mNDCG 30.5 30.9 38.8 34.1 mRR 30.1 31.0 39.6 34.3 Average mAP 27.5 28.5 21.5 37.4 mNDCG 29.4 29.1 22.4 38.0 mRR 34.6 32.3 28.9 40.7
[…]
figurePerformance-abstention curves for the ember-v1 model on the StackOverflow dataset. All methods show increasing curves, indicating effective abstention. The reference-based method, $u_{\text{lin}}$, stands out above the others, demonstrating superior abstention performance. Note that $u^{*}$ and $\tilde{u}$ respectively denote the oracle and the random baseline, associated with the $\rho^{*}$ and $\tilde{\rho}$ mechanisms (Sections 2.2 and 4.3).
We report nAUC (Equation 8) for all methods described in Section 3, averaged model-wise (Table 5.1), and illustrate the dynamics of the performance-abstention curves for the ember-v1 model on the StackOverflow dataset in Figure 5.1.
@@
Reference-based abstention works better. The value of reference-based abstention approaches is demonstrated in Table 5.1, in which we notice that the linear-regression-based method $u_{\text{lin}}$ performs on average largely better than all reference-free baselines, edging out the best baseline strategy ($u_{\text{std}}$) by almost 10 points on the mAP metric.
Having shown the superiority of reference-based abstention methods across the entire benchmark, we conduct (unless otherwise specified) the following experiments on the ember-v1 model using mAP for evaluation (main metric in the MTEB leaderboard (Muennighoff et al., 2022)), and compare $u_{\text{lin}}$ to $u_{\text{std}}$, respectively the best reference-based and reference-free methods mAP-wise.
@@
[…]
figureAbstention performance (nAUC in %) vs. no-abstention performance (mAP), for all models and datasets, using $u_{\text{lin}}$ as a confidence function. Each data point represents a model-dataset pair.
[…]
tablePearson’s correlations between no-abstention mAP and abstention nAUC, for each dataset. General correlation is computed over the whole scatter plot, i.e., across all dataset-model pairs. Dataset Correlation SciDocs 0.85 AskUbuntu 0.38 StackOverflow 0.90 Alloprof 0.56 CMedQAv1 0.94 Mmarco 0.63 General 0.73
We investigate whether a correlation exists between abstention effectiveness and model raw performance (i.e., without abstention) on the reranking task. For each model-dataset pair, we compute both no-abstention mAP and abstention nAUC, using $u_{\text{lin}}$ as a confidence function (Figure 5.2, Table 5.2).
A better ranker implies better abstention. Figure 5.2 and Table 5.2 showcase a clear positive correlation between model raw performance and abstention effectiveness, albeit with disparities from one dataset to another. Better reranking methods naturally produce more calibrated score distributions, which in turn leads to improved abstention performance. This highlights the value of incorporating abstention mechanisms into high-performing retrieval systems and reinforces that their purpose is not to turn weak rankers into strong ones, but to enhance the quality of models that already perform reasonably well. For thoroughness, we evaluated a range of open-source models with varying performance levels, though in practice, users typically select models they already know to be sufficiently effective.
@@ 5.3 Threshold Calibration
[…]
tableMAEs for target vs. achieved abstention rates and performance levels (in %) (model: ember-v1, dataset: StackOverflow, metric: mAP). Target Abstention 10% 50% 90% Rate $u_{\text{std}}$ 1.10 1.80 1.08 $u_{\text{lin}}$ 1.07 1.79 1.09 mAP $u_{\text{std}}$ 1.24 1.52 3.09 $u_{\text{lin}}$ 1.22 1.33 1.51
[…]
tablenAUCs averaged test-set-wise (model: ember-v1, metric: mAP). Method $u_{\text{max}}$ $u_{\text{std}}$ $u_{\text{1-2}}$ $u_{\text{lin}}$ Reference set Alloprof 30.1 32.7 17.1 31.4 AskUbuntu 28.5 32.1 19.8 34.2 CMedQAv1 27.9 31.1 17.2 34.2 Mmarco 27.5 32.2 18.9 34.9 SciDocs 23.0 23.5 22.3 11.6 StackOverflow 29.3 30.7 11.9 20.6 Average 27.7 30.4 17.9 27.8
In practical industrial settings, a major challenge with abstention methods lies in choosing the right threshold to ensure a desired abstention rate or performance level on new data. To evaluate threshold calibration quality, we choose three target abstention rates (10%, 50%, and 90%) and rely on the reference set [^9] In this setting, access to a reference set is essential in order to determine which threshold value corresponds to a given abstention rate or level of performance. to infer the corresponding threshold and level of performance (here, the mAP). Then, the Mean Absolute Error (MAE) between target and achieved values on test instances is reported (Table 5.3), averaging across 1000 random seeds on the StackOverflow dataset. [^10] For example, if the mAP corresponding to a target abstention rate of 50% in the reference set is 0.6 and the values achieved on test set on 2 different seeds are 0.5 and 0.7, the MAE is be equal to $1/2\times(|0.6-0.5|+|0.6-0.7|)=0.1$.
$u_{\text{lin}}$ is better calibrated than $u_{\text{std}}$. The lower part of Table 5.3 shows that the MAE between the target and achieved mAP increases with the abstention rate, which is logical because a higher abstention rate reduces the number of evaluated instances and therefore increases volatility. However, we observe that the MAE increases much less significantly for the $u_{\text{lin}}$ method than for the $u_{\text{std}}$ method, with the latter having a MAE twice as high at a target abstention rate of 90%. This provides strong evidence that $u_{\text{lin}}$ is more reliable at high abstention rates.
@@ 5.4 Domain Adaptation Study
In a practical setting, it is rare to have a reference set that perfectly matches the distribution of samples seen at test time. To measure the robustness of our method to data drifts, we fit them on a given dataset and evaluate on all others.
Reference-based approaches are globally not robust to domain changes. From Table 5.3, we see on average that the reference-based method underperforms reference-free baselines when fitted on the "wrong" reference set. However, certain reference datasets seem to enable methods to generalize better overall, even out of distribution, which is mainly an artifact of the number of positive documents per instance (further insights in Appendix C).
@@ 5.5 Reference Set Minimal Size
[…]
tableMinimum reference set size for $u_{\text{lin}}$ to outperform $u_{\text{std}}$ (model: ember-v1, metric: mAP). Dataset Break-Even SciDocs 12 AskUbuntu 9 StackOverflow 19 Alloprof 58 CMedQAv1 50 Mmarco 80 Average 38
[…]
figurenAUC vs. reference set size for $u_{\text{lin}}$ and $u_{\text{std}}$ on all datasets (model: ember-v1, metric: mAP).
@@
Small reference sets suffice. From Figure 5.5 and Table 5.5, we note that $u_{\text{lin}}$ requires only a fairly small amount of reference data (less than 40 samples on average) to outperform $u_{\text{std}}$, which demonstrates the effectiveness of using a reference set even when data is scarce. This provides strong evidence on the applicability of our method to practical industrial contexts, the labeling of around 50 samples appearing perfectly tractable in such settings.